An ideal gas expands isothermally ( T = 410 K) from a volume of 2.50 l and a

pressure of 7.5 atm to a pressure of 1.0 atm. What is the entropy change for this

process? [Ans = 9.3 J/K]


1
Expert's answer
2022-01-18T13:42:55-0500

Find the final volume:


"V_2=V_1\dfrac{P_1}{P_2}=18.75\text{ L}."


The entropy change for this process:


"\Delta S=R\ln\dfrac{V_2}{V_1}=16.8\text{ J/K}."